# OmniQuery: Contextually Augmenting Captured Multimodal Memory to Enable Personal Question Answering

### v1

@@ 9.2. Enriching Memory Data and Visual Intelligence
Integrating additional data sources: Personal knowledge extends beyond photo albums and exists across various applications. While our participants’ photo albums included screenshots of emails, calendar events, and chat histories, these represent only a fraction of the broader personal information available in other communication and social interaction apps. Incorporating data from such sources could significantly enhance OmniQuery’s contextual understanding, allowing for more complex queries and richer memory retrieval. However, integrating these additional data sources presents substantial privacy and ethical challenges. While our evaluations were conducted entirely on users’ local machines and did not explore privacy-preserving implementations in detail, existing research efforts, such as those focused on differential privacy and on-device machine learning, offer promising directions for secure and privacy-aware deployment. Additionally, commercial tools like Apple Intelligence’s private cloud computing serve as examples of ongoing progress in protecting user data while enabling advanced memory retrieval capabilities.
@@
Enhancing visual intelligence: As discussed in Section 8.6.2, questions related to social interactions remain challenging due to the current lack of advanced features like facial recognition for person identification. Future iterations of OmniQuery could integrate such capabilities (with appropriate user consent), enabling the system to track individuals across various memories. This enhancement would support new use cases, such as monitoring social patterns or tracking progress over time, significantly improving the system’s capacity for memory augmentation and retrieval. Additionally, we propose exploring the design and implementation of a comprehensive taxonomy of personal knowledge domains. This would allow users to selectively activate or deactivate specific domains, such as enabling “Social Interactions and Relationships” to infer personal connections while disabling “Personally Identifiable Information” to prevent the system from processing sensitive data like IDs or SSN numbers in photos. This modular approach could enhance both user control and privacy.
@@
Augmenting with future AR technologies: A limitation of personal memory capture is the potential for missed moments when users either forget or are unable to document an experience. As AR technology advances, OmniQuery’s memory augmentation and retrieval capabilities could be seamlessly integrated into AR systems, allowing for more passive and context-aware memory capture. AR devices could leverage real-time contextual triggers (Li et al., 2024b) to proactively surface relevant memories or information, offering proactive assistance in pervasive AR environments. This integration would enhance the user experience by making memory retrieval more intuitive and contextually relevant. However, such passive data capture raises even more significant privacy concerns, which will require future research into secure, privacy-preserving implementations to ensure the responsible use of AI in these settings.
@@ 9.3. Preserving Privacy
As was discussed in above subsections, protecting users’ privacy is crucial in developing future personal AI assistants, including but not limited to handling personal data such as media in albums, chat histories or browsing history. Users have limited control over how their data is handled and must rely on service providers’ adherence to privacy protocols. In this subsection, we take a step further to discuss more robust and rigorous measures should be adopted in real-world settings, where the amount of personal data is huge, making approaches like manual filtering in OmniQuery’s evaluation infeasible.
One way is to incorporate more advanced data protection techniques on cloud servers including data anonymization (Majeed and Lee, 2021) and encryption (Nadeem and Javed, 2005), while preserving the computational capabilities of large models via online computing. The other approach is leveraging on-device computing, where all data processing occurs locally on the user’s device, ensuring full control over users’ own data. Recent advances in model compression (Hohman et al., 2023) have made it possible to run large model on smaller devices like smartphones. As OmniQuery is designed to be model-agnostic, it is able to work with different model sizes. While smaller, compressed on-device model may result in reduced performance, future work should focus on developing curated datasets and benchmarks to rigorously evaluate OmniQuery ’s performance across different model sizes (e.g., LLaMAs (Touvron et al., 2023) and Phi-3 (Abdin et al., 2024)). This would provide a deeper understanding of how model size impacts both privacy and system effectiveness. concern is that users do not have direct control over their own data when leveraging powerful AI models via online servers.
@@ 10. CONCLUSION
In this paper, we present OmniQuery, which enhances personal question answering on captured memories. To inform the design of OmniQuery, we ran a one-month diary study to collect and analyze realistic user queries, which were then used to generate a taxonomy of contextual information, including atomic context, composite context and semantic knowledge. The taxonomy guided the design of an augmentation pipeline of captured memories, which involves structuring individual captured memories, and processing multiple captured memories using sliding windows to identify composite contexts and semantic knowledge.
We then developed a personal question-answering system, which first augment the user query, then retrieve the related captured memories, and finally using the retrieved memories to generate comprehensive answers using an LLM. We evaluated OmniQuery against a baseline system with a user evaluation of 10 participants who tested 137 queries in total. The results show that OmniQuery is effective in answering users’ queries with an accuracy of 71.5%, and winning or tying 74.5% of the time in direct comparison with the baseline. We also performed qualitative analysis on participants’ reaction and feedback during the evaluation, with the findings demonstrating potential of OmniQuery and providing valuable insights into possible enhancements for future systems alike.
@@ A.2. Inferring Semantic Knowledge
A.3. Generating Answers Based on Retrieved Results
@@ Appendix B Baseline Implementation
While there is no already-existing system designed for answering personal questions on captured memories, we manually designed and implemented a system as the baseline to comare with OmniQuery.
Similar to OmniQuery, the baseline system also adopts a RAG architecture to adapt to the large number of captured memories. We utilized the basic structure of RAG illustrated in (Gao et al., 2023), which involves (1) indexing the external data sources with embedding models, (2) leverage vector-based search to retrieve the top K relevant data instances (3) based-on the retrieved data, utilizing a powerful LLM to generate the final answer. Note that typical RAG systems require a chunking phase, where long documents are split into smaller chunks for more precise matching and retrieval of relevant information. In our case, each captured memory already represents a limited amount of information and is naturally separated. Therefore, we treat each captured memory as an individual chunk.

### v2

@@ 9.3. Enriching Memory Data and Visual Intelligence
Integrating additional data structure and sources. Personal knowledge extends beyond photo albums and exists across various applications. While our participants’ photo albums included screenshots of emails, calendar events, and chat histories, these represent only a fraction of the broader personal information available in other communication and social interaction apps. Incorporating data from such sources could significantly enhance OmniQuery’s contextual understanding, allowing for more complex queries and richer memory retrieval. While OmniQuery already extracts hierarchically structured information from raw photo album data in the form of atomic and composite contexts, future work can explore using other explicit data structures such as graphs and trees to organize data from multiple sources. Integrating these additional data sources also presents substantial privacy and ethical challenges. While our evaluations were conducted entirely on users’ local machines and did not explore privacy-preserving implementations in detail, existing research efforts, such as those focused on differential privacy and on-device machine learning, offer promising directions for secure and privacy-aware deployment. Additionally, commercial tools like Apple Intelligence’s private cloud computing serve as examples of ongoing progress in protecting user data while enabling advanced memory retrieval.
@@
Enhancing visual intelligence. Queries related to social interactions remain challenging due to the current lack of advanced features like facial recognition for person identification. Future iterations of OmniQuery could integrate such capabilities (with appropriate user consent), enabling the system to track individuals across various memories. This enhancement would support new use cases, such as monitoring social patterns or tracking progress over time, significantly improving the system’s capacity for memory augmentation and retrieval. Additionally, we propose exploring the design and implementation of a comprehensive taxonomy of personal knowledge domains. This would allow users to selectively activate specific domains, such as enabling “Social Interactions and Relationships” to infer personal connections while disabling “Personally Identifiable Information” to prevent the system from processing sensitive data like IDs or SSNs in photos. This modular approach could enhance both user control and privacy.
@@
Augmenting with future AR technologies. A limitation of personal memory capture is the potential for missed moments when users either forget or are unable to document an experience. As AR technology advances, OmniQuery’s memory augmentation and retrieval capabilities could be seamlessly integrated into AR systems, allowing for more passive and context-aware memory capture. AR devices could leverage real-time contextual triggers (Li et al., 2024b) to proactively surface relevant memories or information, offering proactive assistance in pervasive AR environments. This integration would enhance the user experience by making memory retrieval more intuitive and contextually relevant. However, such passive data capture raises even more significant privacy concerns, which will require future research into secure, privacy-preserving implementations to ensure the responsible use of AI in these settings.
@@ 9.4. Preserving Privacy
As discussed above, protecting users’ privacy is crucial in developing future personal AI assistants, including but not limited to handling personal data such as media in albums and chat and browsing histories. Users have limited control over how their data is handled and must rely on service providers’ adherence to privacy protocols. In this subsection, we take a step further to discuss more robust and rigorous measures that should be adopted in real-world settings, where the immense amount of personal data makes approaches like manual filtering in OmniQuery’s evaluation infeasible.
One way is to incorporate more advanced data protection techniques, such as data anonymization (Majeed and Lee, 2021) and encryption (Nadeem and Javed, 2005), while preserving the computational capabilities of large models via online computing. The other approach is leveraging on-device computing, where all data processing occurs locally on the user’s device, ensuring full control over users’ own data. Recent advances in model compression (Hohman et al., 2023) have made it possible to run large model on smaller devices like smartphones. As OmniQuery is designed to be model-agnostic, it is able to work with different model sizes. While smaller, compressed on-device model may result in reduced performance, future work should focus on developing curated datasets and benchmarks to rigorously evaluate OmniQuery ’s performance across different model sizes (e.g., LLaMAs (Touvron et al., 2023) and Phi-3 (Abdin et al., 2024)). This would provide a deeper understanding of how model size impacts privacy and system effectiveness.
@@ 10. CONCLUSION
We present OmniQuery, a pipeline that enhances personal question answering on captured multimodal memories. Informed by an one-month diary study, OmniQuery’s design responds to real-world user queries and synthesizes a contextual taxonomy of captured memories. Our pipeline design of structuring individual captured memories and identifying composite context and semantic knowledge using a sliding window technique was used to develop a question-answering system, which outperformed a baseline RAG system in both perceived accuracy and completeness. Unlike existing research and commercial tools focused on intelligent image retrieval in smartphone albums, OmniQuery is the first to tackle complex and nuanced personal queries, moving beyond simple object or information piece retrieval. With further attention to privacy-preserving measures, we believe OmniQuery holds significant potential to evolve into a comprehensive multimodal interactive memory assistant, empowering users to revisit, engage with, and manage their personal memories with greater depth and control.
@@ A.2. Inferring Semantic Knowledge
A.3. Query Augmentation
System instruction:
augment_query = Given a query and todays date, identify the contextual filters. Contextual filters may include:
temporal information: e.g., "last week."
location information: e.g., "Hawaii."
visible objects: e.g., "poke bowl."
people Seen: e.g., "people at the conference."
activities performed: e.g., "ordering in a restaurant."
and more complex contexts such as events or travel which consists of multiple atomic contexts mentioned above: e.g., "traveling to Hawaii." (activity: travel, location: Hawaii).
The query may not contain detailed contextual filters. In such cases, make reasonable inferences. For example, for query "What products did I buy from Sephora", the result could be obtained from a Sephora receipt. Thus inferred contextual filters for objects might be "makeup/skincare products or receipts."
Output a JSON object with the key augmented_query’, including the sub-keys start_date’, end_date’, location’, objects’, people’, activities’, and complex_context’. Each sub-key should be a single string. Leave any sub-key empty if not applicable.
A.4. Generating Answers Based on Retrieved Results
@@ Appendix B Baseline Implementation
While there is no already-existing system designed for answering personal questions on captured memories, we manually designed and implemented a system as the baseline. Similar to OmniQuery, the baseline system also adopts an RAG architecture to adapt to the large number of captured memories. We utilized the basic structure of RAG illustrated in (Gao et al., 2023), which involves (1) indexing the external data sources with embedding models, (2) leverage vector-based search to retrieve the top K relevant data instances (3) based-on the retrieved data, utilizing a powerful LLM to generate the final answer. Note that typical RAG systems require a chunking phase, where long documents are split into smaller chunks for more precise matching and retrieval of relevant information. In our case, each captured memory already represents a limited amount of information and is naturally separated. Therefore, we treat each captured memory as an individual chunk.
Figure 1. Structure of the baseline implementation.
@@
Appendix C Failure Cases Analysis
C.1. Failure Case Categorization
Analyzing failure cases is important to understanding the limitations and improving the design of OmniQuery. Among the 137 queries tested in the study, we identified 25 queries with inaccurate results (OmniQuery’s UPA $\leq$ 2) as failure cases. Additionally, we reached out to participants asking them to manually retrieve the correct memories for these failures. Through this analysis, we categorized them and propose future solutions for each:
Case 1: Ambiguity (8 cases): Ambiguity in language-based interaction was the cause of failure in certain cases based on our analysis. Specifically, such ambiguities can be categorized as follows:
  1. (1)
    Wording ambiguity: P3 asked ”what was the pool place I went in NYC”. While they were referring to a billiards place (less ambiguous term), OmniQuery interpreted it as a swimming pool, resulting in retrieval failure (Figure 2a).
  2. (2)
    Reference ambiguity due to multiple valid answers: Some queries have multiple potential answers. For example, P4 asked ”What is the price of the medicine I bought?” They were referring to the most recent hospital visit, but OmniQuery retrieved a different medicine receipt from the memory, leading to failure in answering users’ question.
  3. (3)
    Contextual ambiguity: P6, a photographer, asked, ”How many times did I work as a photographer in the past few months?” As they also enjoy personal photography, the boundary between photos taken for work and those taken for leisure is ambiguous, causing the system to fail.
The above presents the challenges OmniQuery faces in addressing ambiguity, both in understanding user queries and in analyzing retrieved results. For a detailed discussion on strategies to address these uncertainties, please refer to Section 9.2.
Case 2: Lack of contextual cues (7 cases): Several failure cases occurred due to insufficient contextual information to associate the target memory with the input query. There are two types:
  1. (1)
    Target context being too implicit: As shown in Figure 2b, P4 asked “What is the content of the last meeting with my advisor last week?” The target memory, a photo of a notebook page, lacks contextual cues to associate it with the meeting. This led to retrieval failure. In such cases, OmniQuery should ask the user to clarify or iterate on the query (e.g., specifying the type of memory if the user has an idea).
  2. (2)
    People Identity and Metadata: For example, P5 asked “Where did I travel with two Korean friends last month?” OmniQuery currently lacks access to facial recognition or metadata that can identify and associate individuals by attributes such as race. This also led to retrieval failure when answering this question. Future work could integrate with platforms like Google Photos or Apple Albums, which group photos by individuals using facial recognition. Additionally, users could manually add metadata via linking photos to contacts or descriptive tags (e.g., “friends from summer school”), enabling the system to handle such queries better.
Figure 2. Four exemplar failure cases: (a) lack of context, (b) wording ambiguity, (c) information loss during processing and (d) redundancy-induced failure.
Case 3: Information loss during text-based preprocessing (3 cases): OmniQuery currently adopts a text-based augmentation to extract atomic, composite context and semantic knowledge, which might lead to information loss. For example, P10 asked “What was the brand of the golf shirt I saw in the store?” The brand logo was barely visible in the bottom-left corner in the target memory (Figure 2c) and was not captured during the preprocessing. This led to retrieval failures when answering the question. Future work could integrate current text-based retrieval with advanced multimodal retrieval models (e.g., ColPali (Faysse et al., 2024)), which are capable of keeping more details during the retrieval process.
Case 4: Redundacny-induced failure (3 cases): Similar to the “needle in a haystack” challenge (Li et al., 2024c), OmniQuery’s performance might degrade when too many memories are retrieved during the retrieval phase. For example, P7 frequently takes photos of their cats and when they asked “I took a picture of my cat on the desk with a red hat. Can you find it?” OmniQuery tries to retrieve all relevant memories about their cats, resulting in failure to find the correct one (Figure 2d). In contrast, as the baseline system always retrieves a fixed number of results, it is able to identify the correct answer by narrowing down the search space. To address this, query-aware filtering process such as reranking (Dong et al., 2024) could be employed to narrow down the search space. Additionally, employing a Top-K retrieval strategy (Khattab et al., 2020) could provide users with more options to enhance overall performance.
Case 5: Subjectivity-induced failure (1 case): P2 asked, “Give me the best selfie I took.” However, the subjective nature of “the best” made it difficult for OmniQuery to determine the correct answer. To address such cases, future systems could integrate user preferences (e.g., leveraging marked favorites) or enable interactive clarifications (e.g., asking, “Do you prefer an indoor or outdoor selfie?”) to better align with user intent.
Case 6: Target memory out of scope (3 cases): In three cases, the target memory was unavailable. In one case, the target memory was outside the 100-file range while the participants thought it had been included. In the other two cases, participants mistakenly thought they had captured the memory, but it was not actually captured. It is important to communicate such uncertainties to users when the system believes a query to be not answerable. For detailed discussion, please refer to Sec 9.2.
C.2. Perceived Feelings from Participants
We also present cases when participants reacted negatively to the answers. All participants encountered cases where the answers are inaccurate. Some were incomplete (e.g., P1 believed that they visited mroe than a few churches on the trip to Barcelona, but answers provided only two of them). Some were presumptive (e.g., P7 asked about recent social events, where the answers gave a piece of memory on a museum visit and explained that visiting museums is “likely with other people.” However, P7 visited the museum alone). Some were making mistakes (e.g., P7 asked for the mostly visited attractions but both system mistakenly answered a museum, which was because P7 took a lot of museum pictures and both systems failed to recognize that they were the same visit.) Some even more challenging questions that caused failure of both systems include questions relate to a specific person. For example, P8 asked about her significant other and P5 asked about their “Korean friend” met in a trip. These cases represent the difficulty of understanding the nuances of personal relationships with personal album data.
C.3. Cases Where the Baseline RAG Outperforms OmniQuery
As discussed above, in cases where there is redundancy in the retrieved results before generating the answer, the baseline RAG system may perform better than OmniQuery because it retrieves a fixed number of memories, narrowing the input context and reducing noise during answer generation. To further understand the comparison, we also examined cases where the baseline performs better even when both results are reasonably accurate (UPA $\geq$ 3). Typically, in cases where there is ambiguity in the query, while OmniQuery might provide a relatively accurate result, the baseline often produces a more binary outcome (either highly accurate or highly inaccurate) due to its narrower retrieval scope. This reduces ambiguity but also limits its ability to handle complex contexts.
Appendix D Curating a Fixed Dataset Discussion
OmniQuery is evaluated through real user data, and its effectiveness can be further evaluated on a fixed benchmark dataset. However, curating a fixed benchmark dataset for personal data presents significant challenges:
  • Difficulties in collection of long-term personal data while preserving privacy: Collecting diverse, long-term personal data while preserving participants’ privacy with proper consent and redaction is complex. Prior work like Ego4D ensures privacy by obtaining consent for controlled indoor environments or by de-identifying data through redaction of visible and audible PII (Grauman et al., 2021). However, personal captured memories inherently include interactions with various people and sensitive personal content (e.g., photos of IDs or financial documents), which makes it impractical to obtain universal consent or redact all PII without compromising the data’s utility. A promising solution is advanced generative content replacement (e.g., Xu et al. (Xu et al., 2024)), which replaces sensitive PII with synthetic content, ensuring privacy while preserving the cabality of benchmarking.
  • Difficulties in generating objective and unbiased QA pairs: Personal captured memories and corresponding questions are by definition subjective and sometimes ambiguous. Subjectivity varies across question types. Some focus on objective facts (e.g., “What is the Wifi password?”), while others could be highly subjective (e.g., “What sports were my favorite last year?”). To address this, research methods can be applied to reduce ambiguity and bias, such as leveraging crowdsourcing to create QA pairs from third-person perspectives or assess the objectivity of existing pairs. Future work should aim to develop a taxonomy of question types and explore strategies for guiding crowd workers to assess objectivity or generate improved QA pairs based on the taxonomy.